OpenStudy (anonymous):

6(x - 1) = 4(x - 2) is it -1?

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

yeeeeaaa

OpenStudy (anonymous):

2x = -2 x = -1 right

OpenStudy (anonymous):

6x - 6 = 4x - 8 6x - 4x = -8 +6 2x = -2 x = -2/2 = -1 so correct